# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v3

@@ 3.3 Advantages
Our approach is agnostic to the choice of network architecture for the generative model: it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models [song2020score]. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [xiang2023denoising] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
@@
PRG leverages the infinite-layer structure of continuous-time stochastic flow models [Chen2018], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like SD3 [rombach2022high] and DALL·E 3 [betker2023improving]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering excellent performance with enhanced capacity.
@@
Features extracted by reversing the generative process remain stable along the reverse trajectory, allowing feature use from any point $t\in[0,1]$ for downstream tasks, not just $x_{1}$. These features are robust to various discretization schemes (e.g., Euler, Runge-Kutta) and time steps ($10^{3}$ to $10^{1}$). Our method generalizes well across datasets, enabling fine-tuning of pretrained models for new tasks. Notably, it applies to all diffusion and flow models and has been successfully transferred to community-developed text-to-image models in LABEL:sec:trans.
@@ 4.1 Settings
For generative model selection, we adopt three diffusion and flow-based models: GVP, ICFM, and OTCFM. For pre-training, we followed the protocol from [nichol2021improved], using the Adam optimizer [kingma2014adam] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [liu2021swin], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. We followed DDAE [xiang2023denoising]’s setting (64 resolution, simple data augmentation, no mixup [zhang2017mixup], no cutmix [yun2019cutmix]) for fair comparison, yet our method achieves SOTA diffusion classifier performance. Further details, including hyperparameters and efficiency statistics for training and inference, are provided in Secs. B.1, B.2 and B.3 .
@@
We select state-of-the-art (SOTA) representation models as baselines to evaluate PRG on CIFAR-10 [Krizhevsky2009LearningML], Tiny-ImageNet [Le2015TinyIV], and ImageNet [Russakovsky2014ImageNetLS]. Baselines include classical discriminative models like WideResNet-$28$-$10$ [zagoruyko2016wide] and ResNeXt-$29$-$16\times 64$[xie2017aggregated], as well as diffusion-based generative models such as GLOW [fetaya2019understanding], Energy Model [duvenaud2020your], SBGC [zimmermann2021score], HybViT [yang2022your], and DDAE [xiang2023denoising]. For out-of-distribution (OOD) evaluation, we compare PRG with PGD [madry2017towards], PLAT [wang2020enresnet], AugMix [hendrycks2019augmix], AutoAug [cubuk2019autoaugment], SBGC, and PDE+ [yuan2024pde] on CIFAR-10-C and Tiny-ImageNet-C. All results are from their original papers.
@@ 4.3.1 Better Pre-training Lead to Better Fine-tuning?
$\mathbb{E}_{p(x_{1},x_{0})}\bigl{[}\log p(x_{0}\mid x_{1})\bigr{]}=\mathcal{I}(X_{0},X_{1})\;-\;\mathcal{H}(X_{0}),$
where $\mathcal{H}(X_{0})$ is the (constant) entropy of the data distribution and $\mathcal{H}(X_{1})$ is the (constant) entropy of the Gaussian distribution. Hence, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ faithfully reflects the mutual information $\mathcal{I}(X_{0},X_{1})$ during pre-training. In contrast, during fine-tuning, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ no longer represents $\mathcal{I}(X_{0},X_{1})$, because $\mathcal{H}(X_{1})$ is no longer the entropy of a Gaussian distribution. Nonetheless, the quantity $\mathcal{I}(X_{0},X_{1})+\mathcal{H}(X_{1})\;-\;\mathcal{H}(X_{0})$ still captures the change in $\mathcal{I}(X_{0},X_{1})$ to a large extent. We compute log-probability density $\log p(x_{0}|x_{1})$ using the adjoint method or Neural ODEs [Chen2018] for computational feasibility:
@@
For efficient computation of $\log p(x_{0}|x_{1})$, we employ the Hutchinson trace estimator [Hutchinson1989, Grathwohl2018]:
@@ 4.3.3 Is the Model a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [he2016deep], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
@@
WideResNet-$28$-$10$ [zagoruyko2016wide] $36$ $96.3$
ResNeXt-$29$-$16*64$d [xie2017aggregated] $68$ $96.4$
@@
GLOW [fetaya2019understanding] N/A $84.0$
Energy model [duvenaud2020your] N/A $92.9$
SBGC [zimmermann2021score] N/A $95.0$
HybViT [yang2022your] $43$ $96.0$
DDAE [xiang2023denoising] $36$ $97.2$
[…]
@@
HybViT [yang2022your] $43$ $56.7$
DDAE [xiang2023denoising] $40$ $69.4$
[…]
@@
Supplementary Material
@@ Appendix A Why PRG is effective?
According to the manifold assumption [murphy2023probabilistic, Chapter 20], data lies on a low-dimensional manifold $\mathcal{M}$ of intrinsic dimension $d^{*}$, which is significantly smaller than the ambient dimension $D$. Although, in theory, intermediate latents encode the same information as the original data, using these latents as inputs for downstream tasks is more meaningful. A well-trained flow model effectively extracts data from the low-dimensional manifold $\mathcal{M}$ and lifts it to the ambient space $\mathbb{R}^{D}$, ensuring that every sample in $\mathbb{R}^{D}$ remains semantically meaningful. In contrast, directly sampling from the original data space does not necessarily preserve semantic coherence. For instance, generating a $64\times 64$ image by sampling from a $64\times 64$ multivariate Gaussian distribution would typically result in a meaningless noise-like image resembling a snowflake pattern.
Moreover, since the flow trajectories of an ordinary differential equation (ODE) do not intersect [Chen2018], points within a given region remain confined, thereby preserving topological relationships throughout the transformation process.
@@ A.2 Model-agnosticity
Our method follows the model-agnostic principle as defined in Appendix D.5 of [song2020score], where different architectures achieve comparable encoding quality through flow matching. This indicates that architectural constraints stem from implementation choices rather than the framework itself.
Notably, while early diffusion models primarily relied on U-Nets for practical stability, recent work [bao2023all] has demonstrated that flow matching can be successfully achieved with minimal architectural requirements, such as shallow skip connections. In this work, we achieve state-of-the-art performance across both U-Net (LABEL:sec:unet) and Transformer (LABEL:sec:trans) architectures, further validating the generality of our approach.
@@ B.1 Training Process Details
Tabs. 2 and 3 list the hyperparameters used for both pre-training and fine-tuning across the CIFAR-10, Tiny-ImageNet, and ImageNet datasets.
@@
[…]
@@ B.2 Evaluation of Training Efficiency
[…]
Table 4: Mean time cost per epoch during the training process. * means that the model is trained on $4*8$ A100 GPUs. $t_{\text{span}}$ represents the total sampling length, while $t_{\text{cutoff}}$ indicates the point along the trajectory where fine-tuning begins. For example, $20/10$ means a trajectory spanning $20$ steps from $x_{0}$ to $x_{1}$, with fine-tuning starting from the midpoint of the trajectory.

### v6

@@ 3.3 Advantages
Our approach is agnostic to the choice of network architecture for the generative model: it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models [58]. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [68] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
@@
PRG leverages the infinite-layer structure of continuous-time stochastic flow models [11], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like SD3 [53] and DALL·E 3 [7]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering excellent performance with enhanced capacity.
@@
Features extracted by reversing the generative process remain stable along the reverse trajectory, allowing feature use from any point $t\in[0,1]$ for downstream tasks, not just $x_{1}$. These features are robust to various discretization schemes (e.g., Euler, Runge-Kutta) and time steps ($10^{3}$ to $10^{1}$). Our method generalizes well across datasets, enabling fine-tuning of pretrained models for new tasks. Notably, it applies to all diffusion and flow models and has been successfully transferred to community-developed text-to-image models in Sec. 4.4.3.
@@ 4.1 Settings
For generative model selection, we adopt three diffusion and flow-based models: GVP, ICFM, and OTCFM. For pre-training, we followed the protocol from [49], using the Adam optimizer [31] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [40], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. We followed DDAE [68]’s setting (64 resolution, simple data augmentation, no mixup [74], no cutmix [72]) for fair comparison, yet our method achieves SOTA diffusion classifier performance. Further details, including hyperparameters and efficiency statistics for training and inference, are provided in Secs. B.1, B.2 and B.3 .
@@
We select state-of-the-art (SOTA) representation models as baselines to evaluate PRG on CIFAR-10 [33], Tiny-ImageNet [34], and ImageNet [54]. Baselines include classical discriminative models like WideResNet-2828-$10$ [73] and ResNeXt-$29$-$16\times 64$[69], as well as diffusion-based generative models such as GLOW [19], Energy Model [18], SBGC [77], HybViT [70], and DDAE [68]. For out-of-distribution (OOD) evaluation, we compare PRG with PGD [44], PLAT [66], AugMix [25], AutoAug [15], SBGC, and PDE+ [71] on CIFAR-10-C and Tiny-ImageNet-C. All results are from their original papers.
@@ 4.3.1 Better Pre-training Lead to Better Fine-tuning?
$\small\mathbb{E}_{p(x_{1},x_{0})}\bigl{[}\log p(x_{0}\mid x_{1})\bigr{]}=\mathcal{I}(X_{0},X_{1})\;-\;\mathcal{H}(X_{0}),$
where $\mathcal{H}(X_{0})$ is the (constant) entropy of the data distribution and $\mathcal{H}(X_{1})$ is the (constant) entropy of the Gaussian distribution. Hence, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ faithfully reflects the mutual information $\mathcal{I}(X_{0},X_{1})$ during pre-training. In contrast, during fine-tuning, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ no longer represents $\mathcal{I}(X_{0},X_{1})$, because $\mathcal{H}(X_{1})$ is no longer the entropy of a Gaussian distribution. Nonetheless, the quantity $\mathcal{I}(X_{0},X_{1})+\mathcal{H}(X_{1})\;-\;\mathcal{H}(X_{0})$ still captures the change in $\mathcal{I}(X_{0},X_{1})$ to a large extent. We compute log-probability density $\log p(x_{0}|x_{1})$ using the adjoint method or Neural ODEs [11] for computational feasibility:
@@
For efficient computation of $\log p(x_{0}|x_{1})$, we employ the Hutchinson trace estimator [29, 22]:
@@ 4.3.3 Is the Model a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [23], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
@@
WideResNet-2828-$10$ [73] $36$ $96.3$
ResNeXt-2929-166416*64d [69] $68$ $96.4$
@@
GLOW [19] N/A $84.0$
Energy model [18] N/A $92.9$
SBGC [77] N/A $95.0$
HybViT [70] $43$ $96.0$
DDAE [68] $36$ $97.2$
Our methods
@@
HybViT [70] $43$ $56.7$
DDAE [68] $40$ $69.4$
Our methods
@@
Method Param. (M) Acc. (%)
Discriminative methods
ViT-L/$16$ ($384^{2}$) [17] $307$ $76.5$
ResNet-$152$ ($224^{2}$) [23] $60$ $77.8$
Swin-B ($224^{2}$) [40] $88$ $83.5$
Generative methods
HybViT ($32^{2}$) [70] $43$ $53.5$
DMSZC-DiTXL2 ($256^{2}$) [35] $338$ $77.5$
iGPT-L ($48^{2}$) [10] $1362$ $72.6$
Our methods
PRG-GVP-onlyPretrain ($64^{2}$) $122$ $20.18$
PRG-GVP-XL ($64^{2}$) $122$ $77.84$
PRG-ICFM-XL ($64^{2}$) $122$ $78.12$
PRG-OTCFM-XL ($64^{2}$) $122$ $78.13$
Table 5: Comparison on the ImageNet dataset with various algorithms. All results are reported from their original papers, the values in parentheses indicate the input image size.
4.4 Main Results
4.4.1 Performance on Image Classification
To comprehensively evaluate our method, we conducted experiments on three image classification datasets, as detailed in Tabs. 3, 4 and 5. Our methods achieved accuracies of 97.59%, 71.12%, and 78.1% on CIFAR-10, Tiny-ImageNet, and ImageNet, respectively, surpassing all existing generative approaches, as well as supervised methods like WideResNet-28[73]. However, these results still trail behind recent transformer-based supervised learning architectures like SwinTransformer [40]. We hypothesize that integrating transformer architectures and leveraging VAEs to handle high input resolutions could further enhance the performance. This avenue will be explored in future work.
4.4.2 Out-of-Distribution Robustness
To address the degradation in out-of-distribution (OOD) due to common image corruptions or adversarial perturbations [77], data augmentations and adversarial training are typically employed. However, recent studies [77, 8, 35] have indicated that generative-based classifiers, without requiring additional data, often exhibit superior OOD robustness. As illustrated in Tab. 6 and Sec. C.3, our method showcases remarkable robustness on these two datasets with only simple data augmentation.
Model CIFAR-10 CIFAR-10-C
Clean Corr Severity All w/ Noise
Adversarial Training
PGD *[44] $93.91$ $83.08$ $\downarrow 10.83$ $82.10$ $\downarrow 11.81$
[…]
Noise Injection
RSE *[38] $95.59$ $77.86$ $\downarrow 17.73$ N/A
ENResNet *[66] $83.33$ $74.34$ $\downarrow 8.99$ N/A
Data Augment
AugMix § [25] $95.83$ $89.09$ $\downarrow 6.74$ $88.71$ $\downarrow 7.12$
[…]
Generative Methods
[…]
Our methods
PRG-GVP-S $97.35$ $91.21$ $\downarrow 6.14$ $92.17$ $\downarrow 5.18$
PRG-ICFM-S $97.59$ $92.13$ $\downarrow 5.46$ $93.07$ $\downarrow 4.52$
PRG-OTCFM-S $97.65$ $92.26$ $\downarrow 5.39$ $93.10$ $\downarrow 4.55$
Table 6: (OOD: Image Corruptions) Comparison on CIFAR-10-C, including Noise, Blur, Weather, and Digital corruptions. Results are referenced from original papers or [71, 77]. * denotes ResNet-18 as the base model, while § indicates ResNeXt-29.
Method Param. (M) Cifar Acc. Tiny Acc.
Classic self-supervised learning methods
MAE (ViT-B/16) [41] $86$ $96.5$ $76.5$
SimCLR Res-50-4x [12] $375$ $98.6$ N/A
Generative methods
DDAE-DiT-XL2 [68] $338$ $98.4$ $77.8$
iGPT-L [10] $1362$ $99.0$ N/A
Our methods
PRG-SiT-XL2 $338$ $98.72$ $78.33$
Table 7: Comparison of models pretrained on ImageNet-1k and transferred (fine-tuning) to CIFAR-10 and Tiny-ImageNet.
4.4.3 Transferring Features
A main goal of the pre-training/fine-tuning paradigm is to learn transferrable features. We hypothesize that advancements like powerful pretrained models in the generative model community can contribute to our method. To validate this, we conducted experiments by transferring the pretrained SiT-XL [43] model to two other datasets. Since SiT provides only class-conditional checkpoints, we adopted an unconditional approach by setting the label parameter to null. Specifically, we fine-tuned the model for 28 epochs on the CIFAR-10 dataset and for 45 epochs on the Tiny-ImageNet dataset. As demonstrated in Tab. 7, our method shows superior transfer learning performance. The results suggest that our algorithm benefits from larger datasets and the latent space of generative models.
4.5 Ablation Studies
In this section, we conduct three ablation studies to further evaluate the applicability of our method. Additional experiments, including analyses of loss type, ODE solver choice, dual-task supervision, and $t_{\text{span}}$, are provided in Appendix C.
4.5.1 Generative Model Type
We explored various generative model types, particularly focusing on path selection. To assess the impact of these paths on our method, we conducted experiments with three model types on the CIFAR-10 and Tiny-ImageNet datasets. Following [39], we quantified the straightness of various continuously differentiable trajectories. Interestingly, while all three model variants achieved comparable performance (Tab. 8), models with higher straightness required fewer pre-training steps to attain equivalent performance levels and demonstrated potential for superior outcomes.
Generative Model Type GVP ICFM OTCFM
CIFAR-10 (Accuracy) $97.35$ $97.59$ $97.65$
First to 9797% (Epoch) $162$ $135$ $128$
Straightness $7.36$ $0.34$ $0.11$
Tiny-ImageNet (Accuracy) $70.98$ $71.12$ $71.33$
First to 7070% (Epoch) $157$ $125$ $106$
Straightness $6.25$ $0.54$ $0.15$
Table 8: (Generative Model Type) Performance of different generative model types on the CIFAR-10 and Tiny-ImageNet datasets.
4.5.2 Effect of $\beta$ on Accuracy
During pre-training, we introduced $\beta$ in Eq. 3 to enhance the mutual information lower bound while balancing the downstream task objective. Table 9 shows PRG scores across datasets for different $\beta$ values.
$\beta$ 1 10 100
Acc. (CIFAR/Tiny) 0.9730.973/$0.712$ 0.9740.974/$0.711$ 0.9650.965/$0.673$
Table 9: ($\beta$ values) Accuracy scores of PRG with varying $\beta$ values on the CIFAR-10 and Tiny-ImageNet datasets.
4.5.3 Scaling Up Network Parameters
We investigated the impact of scaling up U-Net models of varying network parameter sizes [49]. Table 10 provides a comparative analysis following fine-tuning on the CIFAR-10 and Tiny-ImageNet datasets. The findings suggest that increasing the model size results in a modest enhancement of final classification accuracy.
Model Param (M) CIFAR Acc. Tiny Acc.
PRG-ICFM-B $32$ $97.35$ $70.94$
PRG-ICFM-S $42$ $97.59$ $71.12$
PRG-ICFM-L $66$ $97.75$ $71.30$
PRG-ICFM-XL $122$ $97.88$ $71.80$
Table 10: (Scaling Capability) Performance gains of increasing U-Net model sizes on the CIFAR-10 and Tiny-ImageNet datasets.
5 Conclusion
This paper proposes to turn a pretrained continuous-time flow/diffusion generator upside-down: running the model backward produces multi-level features that, after light fine-tuning, serve as an unsupervised representation extractor. We systematically investigate the necessary designs of the two-stage pre-training and fine-tuning paradigm. Theoretical analysis supports the efficacy of the first-stage generative pre-training, while empirical verifications provide insights for designing the fine-tuning pipeline. Leveraging these simple yet effect techniques and findings, our method achieves state-of-the-art performance in image classification tasks using generative models. Additional experiments, such as those on OOD problems and transfer learning scenarios, further validate the robustness and other advantages of our approach. However, there remains room for improvement, including the integration of large foundation pretrained generative models from the open-source community and optimizing the training epoch/speed for faster adaptation to downstream tasks.
Acknowledgments.
This work was supported by the Shanghai Artificial Intelligence Laboratory, in part by the National Key R&D Program of China under Grant No. 2022ZD0119301, and in part by the National Natural Science Foundation of China under Grant No. U21A20485 and No. 62406141. The Work was done during Rongkun Xue’s internship at Shanghai AI Laboratory.
Appendix
@@ Appendix A Why PRG is effective?
According to the manifold assumption [47, Chapter 20], data lies on a low-dimensional manifold $\mathcal{M}$ of intrinsic dimension $d^{*}$, which is significantly smaller than the ambient dimension $D$. Although, in theory, intermediate latents encode the same information as the original data, using these latents as inputs for downstream tasks is more meaningful. A well-trained flow model effectively extracts data from the low-dimensional manifold $\mathcal{M}$ and lifts it to the ambient space $\mathbb{R}^{D}$, ensuring that every sample in $\mathbb{R}^{D}$ remains semantically meaningful. In contrast, directly sampling from the original data space does not necessarily preserve semantic coherence. For instance, generating a $64\times 64$ image by sampling from a $64\times 64$ multivariate Gaussian distribution would typically result in a meaningless noise-like image resembling a snowflake pattern.
Moreover, since the flow trajectories of an ordinary differential equation (ODE) do not intersect [11], points within a given region remain confined, thereby preserving topological relationships throughout the transformation process.
@@ A.2 Model-agnosticity
Our method follows the model-agnostic principle as defined in Appendix D.5 of [58], where different architectures achieve comparable encoding quality through flow matching. This indicates that architectural constraints stem from implementation choices rather than the framework itself.
Notably, while early diffusion models primarily relied on U-Nets for practical stability, recent work [3] has demonstrated that flow matching can be successfully achieved with minimal architectural requirements, such as shallow skip connections. In this work, we achieve state-of-the-art performance across both U-Net (Sec. 4.4.1) and Transformer (Sec. 4.4.3) architectures, further validating the generality of our approach.
@@ B.1 Training Process Details
Min LR $=\text{Min\_LR}_{\text{base}}\times\frac{\text{num processes}\times\text{Batch Size}}{512}$
@@ B.2 Evaluation of Training Efficiency
$20/5$ $49\text{s}$ $307\text{s}$ N/A
$20/10$ $103\text{s}$ $681\text{s}$ $2220\text{s}$
2020 $187\text{s}$ $1437\text{s}$ $3480\text{s}$
$32/16$ $70\text{s} 11$$137\text{s}$ N/A
Table 4: Mean time cost per epoch during the training process. * means that the model is trained on 484*8 A100 GPUs. $t_{\text{span}}$ represents the total sampling length, while $t_{\text{cutoff}}$ indicates the point along the trajectory where fine-tuning begins. For example, $20/10$ means a trajectory spanning 2020 steps from $x_{0}$ to $x_{1}$, with fine-tuning starting from the midpoint of the trajectory.